OpenStudy (anonymous):

Remove the symbols of grouping (i.e., brackets, braces, parentheses)and collect terms: 3(x - y) + {5y - [3x - (2x - y)] + 4(x - y)}

OpenStudy (anonymous):

= 3x-3y + (5y - (3x - 2x + y)) + 4x - 4y = =3x-3y + (5y - (x+y)) + 4x - 4y = =3x-3y + (5y - x - y) + 4x - 4y = =3x-3y + (4y - x) + 4x - 4y = =3x - 3y + 4y - x + 4x - 4y = =6x - 3y

OpenStudy (anonymous):

Would you be willing to help me with the others if I post them please?